OpenStudy (anonymous):

The bacteria in a certain culture double every 7.3 hours. The culture has 7,000 bacteria at the start. How many bacteria will the culture contain after 4 hours ?

OpenStudy (amistre64):

hmm, if we knew the rate, we could determine the rest A=Pe^rt A/P = e^rt ln(A/P) = rt ln(A/P) ------ = r t but since at t=7.3 we have that A = 2P ln(2P/P) ------ = r 7.3

OpenStudy (amistre64):

ergo: \[A=7000e^{\frac{ln(2)}{7.3}t}\] calculate for t=4

OpenStudy (anonymous):

well these are the answer options... A) 8,324 bacteria b) 10,234 bacteria c) 24,801 bacteria d) 7,697 bacteria

OpenStudy (amistre64):

then im sure one of them will fit the bill

OpenStudy (anonymous):

is there any way you can make the equation you just did larger i cant see anything after the 7000

OpenStudy (amistre64):

\[\Huge A=7000e^{\frac{ln(2)}{7.3}t}\] thats about the best I can do to it :)

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

yw, good luck :)

OpenStudy (anonymous):

one last thing .. what does the ln stand for

OpenStudy (anonymous):

natural logarithm, log with base e